# Rapid, strongly magnetized accretion in the zero-net-vertical-flux shearing box

### v1

@@ 3 Theory & Methods
Here $\bm{B}$ is the magnetic field, $\bm{U}=-q\Omega x\bm{\hat{y}}+\bm{u}$ is the total flow velocity (including the background shear $-q\Omega x\bm{\hat{y}}$), $\rho$ is the mass density, $\Omega$ is the local rotation rate, which is the natural time unit for the system, $q\equiv-d\ln\Omega/d\ln R=3/2$ for a Keplerian disk, and $c_{s}$ is the (constant) sound speed. In some simulations we remove the rotational effects (i.e., simulate a gravitating shearing slab of plasma), which amounts to dropping the first two terms on the second line of Eq. 1b, or equivalently, to the limit $\Omega\rightarrow 0$ with $q\Omega=S={\rm const.}$ (while retaining vertical stratification, $-\rho\Omega^{2}z\hat{\bm{z}}$). Some other simulations remove the effects of shear, which involves setting $q=0$. The purpose of these modifications is not to explore physically relevant scenarios, but rather as numerical experiments to probe the important physics.
Although idealized, many studies have shown that the turbulent dynamics induced by Eqs. 1a to 1c are widely varied and dynamically complex, involving the natural generation of long-time coherent large-scale fields and flows, and other complex features. The system also has various undesirable features such as a solutions that depend on the vertical boundary conditions, resolution, and box size (e.g., Simon et al. 2012; Shi et al. 2016; Ryan et al. 2017), as well as the obvious neglect of potentially important global effects (e.g., field-line curvature). For our purposes, it retains two important advantages compared to global simulations: first, it is cheaper computationally than global simulations for similar disk resolutions; second, it provides a much simpler and more controllable testbed for understanding the key physics at play in the system, particularly in steady state. We will see that vertical stratification is absolutely crucial to the physics we describe, meaning that the yet simpler unstratified version of the shearing box would be unsuitable.
@@ 3.1 Static vertical equilibria in the local shearing box
Here $P_{\rm th}\equiv c_{s}^{2}\rho$ is the thermal pressure, $P_{B}\equiv B^{2}/8\pi$ is the magnetic pressure, $\beta(z)\equiv P_{\rm th}/P_{B}$, and $H_{\rm th}^{2}\equiv 2c_{s}^{2}/\Omega^{2}$ is the thermal scale height. We will find that the azimuthal field dominates strongly over the other components, implying that so long as the turbulence is modestly subsonic where $\beta\gtrsim 1$ (such that the thermal pressure dominates the turbulent pressure) and modestly sub-Alfvénic where $\beta\lesssim 1$ (such that the mean magnetic pressure dominates its fluctuating part), Eq. 2 provides a relation between $\rho$ and $B_{y}\approx\sqrt{8\pi P_{B}}$. We will also see that the mean vertical outflows that develop do not provide significant contributions to the overall pressure balance, further justifying the magneto-static assumption.
@@ 3.2 Boundary conditions
In all simulations we use periodic boundary conditions in $y$, and “shear-periodic” boundary conditions in $x$, whereby quantities are periodic in the frame that remains stationary on the background shear flow $-q\Omega x\bm{\hat{y}}$, as required for the shearing-box method itself. However, previous works have shown significant dependence of results on the vertical boundary conditions (e.g., Johansen & Levin, 2008; Lesur et al., 2013; Simon et al., 2012; Salvesen et al., 2016a), so we use a number of different choices through this work. All boundary conditions set $u_{x}$, $u_{y}$, and $B_{z}$ to have zero gradient at the vertical boundary, and set the vertical flow $u_{z}$ to have zero gradient unless there is a inflow ($u_{z}z<0$), in which case it is set to $u_{z}=0$ (Stone et al., 1996). The conditions on $\rho$, $B_{x}$, and $B_{y}$ are as follows, with the naming scheme used to label simulations in table 1:
@@
Zero-gradient boundary conditions for $B_{x}$ and $B_{y}$, while $\rho$ is extrapolated into the ghost cells assuming the initial hydrostatic equilibrium as $\rho(z)=\rho_{e}\exp(-z^{2}/H_{\beta}^{2})/\exp(-z_{e}^{2}/H_{\beta}^{2})$ (where the $z_{e}$ and $\rho_{e}$ here refer to the values at the boundary). Following Simon et al. (2011), these boundary conditions have been used commonly as a simple alternative to periodic boundary conditions that allow the total toroidal and radial magnetic flux in the system to change.
@@
Following Simon et al. (2013) and Salvesen et al. (2016a), these are the same as Outflow, except that they also extrapolate $B_{x}$ and $B_{y}$ using the same scheme as the density, as $B_{i}(z)=B_{i,e}\exp(-z^{2}/H_{\beta}^{2})/\exp(-z_{e}^{2}/H_{\beta}^{2})$. This enhances the removal of magnetic flux from the domain.
@@
Outflow, Simon13, and Power-law effectively differ only in the gradient of $\bm{B}_{\perp}$ and $\rho$ assigned at the boundary, as set by the various extrapolation techniques. For context, the Simon13 hydrostatic scheme (Gaussian extrapolation) is most extreme, fixing quantities to larger negative gradients; the power-law scheme is more moderate; while outflow conditions (with zero gradient) are obviously least extreme. From a close examination of the shape of $\rho$ near the boundaries with Outflow or Simon13, it is clear that the latter forces the profile to be steeper than is natural. This was the primary motivation for the new Power-law scheme — to disturb the solution as little as possible from the profiles that are naturally set up in coronal regions above the midplane.
@@ 3.3 Initial conditions and numerical setup
Following Salvesen et al. (2016a, b), most of our simulations are initialized with the magnetized constant-$\beta$ equilibrium (3). Most simulations use a vertical domain size $z\in(-L_{z}/2,L_{z}/2)$ with $H_{\beta}^{2}=2$ and $L_{z}=10$, giving significant space away from the midplane to form a low-$\beta$ corona (it transpires that this is necessary to maintain the low-$\beta$ state). Based on the results of Simon et al. (2013) and subsequent works, we use $L_{x}=L_{z}$ and $L_{y}=2L_{z}$ for most runs, but also test the effect of increasing and decreasing both $L_{x}$ and $L_{y}$ by a factor of two (App. A). The equilibrium is then defined by its initial $\beta=\beta_{y0}=8\pi P_{\rm th}/B_{y}^{2}$, which sets the simulation’s isothermal sound speed as $c_{s}^{2}=\Omega^{2}H_{\beta}^{2}(1+\beta^{-1})^{-1}$. In table 1, we label simulations by this initial $\beta$, although the midplane $\beta$ invariably decreases significantly from this initial value as the field redistributes and flux is lost through the boundaries.
@@
Nearly all simulations use the piecewise parabolic method with the Van-Leer time integrator and the basic HLL Reimman solver. Although this solver is more diffusive than the HLLD solver, we have found this low-$\beta$ turbulence to be rather sensitive to numerical instabilities, also requiring very conservative choices for the CFL timestep criterion. The strongly magnetized simulations are also quite expensive due to the small timesteps required to capture the fast Alfvén speeds in the very low-$\beta$ coronae. To help mitigate this, and given that highly supersonic turbulence drives large density contrasts in these low-$\rho$ regions, we also apply a density floor of $\rho_{\rm flr}=10^{-4}$ of the initial density maximum in most simulations, as used in previous works (e.g., Salvesen et al. 2016a; another approach is to limit the Alfvén speed Johansen & Levin, 2008). Unfortunately, this density floor was found to have adversely affected some simulations by limiting the mean density profile in the strongly magnetized corona. The issue occurs in the high-$\beta$ states at lower $H_{\rm th}$ (see below), which are not our main focus in this work, because the thermal density profile leads to an extremely sharp drop off away from the midplane (unlike the low-$\beta$ states, which develop a much flatter power-law profile). In App. C we explicitly test the effect of changing the density floor on the low-$\beta$ state, observing only tiny changes to mean profiles or transport rates.
In all simulations except those with net vertical flux, the mass loss over the duration of the simulation is minimal. In the the strongest NVF case with $\beta_{z}=100$, we implement a mass replenishment by multiplying the density by a factor chosen to maintain constant total mass at every time step (as done in previous work; Bai & Stone, 2013).
Note that, compared to previous works, we have run most simulations for a relatively short time (in units of $\Omega^{-1}$). This is motivated by a desire to redeploy computational resources to explore higher resolutions (as needed for $\beta\ll 1$ and small $H_{\rm th}$) and different simulations, rather than measuring precise values of $\alpha$ via long-time statistics. In support of this choice, we note that turbulence in stratified shearing boxes does not yet seem to be converged even at exceptionally high resolution (Ryan et al., 2017), and detailed study along such lines is beyond the scope of this work. To ensure that the system is truly converged in time, we have run one simulation of the low-$\beta$ state to beyond $100$ orbits; this revealed a nearly constant statistical steady state with only small changes over long timescales.
@@ 3.4 Notation & Units
The shearing-box system is normalized to the arbitrary lengthscale $L_{0}$, so there is freedom in defining the simulation’s units. In order to more clearly compare cases with strong magnetization, unlike most previous works, we do not normalize our box lengthscales to the thermal scale height $H_{\rm th}$. Instead, in most simulations, because $H_{\beta}^{2}=2$ and $\beta<1$, we have $H_{\rm th}<1$ in box units. However, everything in the system could, if desired, be rescaled to $H_{\rm th}$, in which case our domain sizes would vary significantly between simulations and be much larger than $(L_{x},L_{y},L_{z})=(10,20,10)$ (this would also rescale the density and velocity). Indeed, we find that small $H_{\rm th}$ (in box units) seems to be required to see the low-$\beta$ state, which is equivalent to saying that very large domains $L_{z}\gtrsim 20$ are needed in thermal units. All plots and diagnostics are normalized to be dimensionless so can be meaningfully compared to previous works, showing good order-of-magnitude agreement where appropriate (see table 1).
@@ 4 The Parker and magnetorotational instabilities
Here we provide an overview of the magnetorotational and Parker instabilities (Parker, 1958; Balbus & Hawley, 1991), which seem to be the linear instabilities that drive the dynamics. While these instabilities and their relation to each other have been extensively explored in the literature (e.g., Foglizzo & Tagger, 1995; Balbus & Hawley, 1998) we focus here on the less-well-studied regime with $\beta\lesssim 1$ trans- or super-thermal magnetic fields, so it is helpful to collect and simplify some past results for more straightforward application to our simulations.
@@ 4.1 Magnetorotational instability
Also worth mentioning are other related, but fundamentally global instabilities at $\beta\ll 1$, which were first characterized by Pessah & Psaltis (2005) (the relevant dispersion relation was derived but left unexplored in Chandrasekhar 1961; see also Das et al. 2018). These instabilities are current driven, depending on the radial gradient of the (super-thermal) toroidal field,
@@ 4.2 Parker instability
The Parker instability (Parker, 1958) is driven by the vertical stratification of the magnetic field. Most commonly, it is studied while ignoring the effect of the radial shear; indeed, Shu (1974) argued that the instability cannot ever be fully stabilized by shear or rotation, because the small-scale radial modes remain unstable (although the growth rates of longer-wavelength modes can be suppressed and shear can play an important role by changing the radial wavenumber in time; Foglizzo & Tagger, 1994, 1995). Convenient expressions are found in Goedbloed & Poedts (2004) Chapter 7, who consider a gravitating slab threaded by a perpendicular magnetic field $B_{y}(z)$ (without shear or rotation). Because the background equilibrium and gravitational force $\hat{g}=-\Omega^{2}z$ are $z$ dependent, a correct treatment requires solving the boundary-value differential equation in $z$; however, useful insight can be gained by considering local modes that vary rapidly over the length scale $L\sim z\sim B_{y}/B_{y}^{\prime}$ on which the background varies (i.e., $\delta f\sim\exp(i\kappa z)$ with $\kappa L\gg 1$). This yields an effective local dispersion relation for $\omega(k_{x},k_{y},\kappa)$, which shows that two different types of modes can dominate the instability in different regimes: (i) the “pure interchange mode” with $k_{\|}=k_{y}=0$, which does not bend the magnetic field, and (ii) the “quasi-interchange mode,” with $k_{x}\gg k_{y}$ but $k_{y}\neq 0$, which bends the field. We obtain simple expressions below by simplifying those of Goedbloed & Poedts (2004) assuming an isothermal equation of state, a $\beta\ll 1$ azimuthally directed mean field, and linear vertical gravity $\hat{g}=-\Omega^{2}z$ [^2] In the notation of Goedbloed & Poedts (2004), the disk atmosphere of interest becomes $\Gamma=-\hat{g}\,d\ln\rho/dz$, $\Gamma_{B}=c_{s}^{-2}\hat{g}^{2}$, $\Gamma_{0}=c_{s}^{-2}\hat{g}^{2}/(1+2/\beta)^{2}$, with $X^{2}\equiv\Gamma/\Gamma_{B}$. The quasi-interchange is dominates for $\Gamma_{0}\leq\Gamma<\Gamma_{B}$ and the pure interchange dominates for $\Gamma\leq\Gamma_{0}$.
@@
Figure 1: Space-time ‘butterfly’ plots for a range of ZNVF simulations with different initial conditions and sound speeds. Each case shows $\overline{B}_{y}(z,t)$ in the top subpanel and $\overline{B}_{x}(z,t)$ in the bottom subpanel. In the top row, we see cases that transition into the high-$\beta$ weakly accreting state ($\beta$1, $\beta$0.3, and $\beta$0.1) because they do not retain sufficient magnetic flux during their initial relaxation phase; these develop the well-studied dynamo cycles. The bottom panels show various cases that maintain the lower-$\beta$ strongly accreting state ($\beta$R0.1, $\beta$0.05, and $\beta$0.01); these maintain a quasi-steady state with $\beta\lesssim 1$ in the midplane, losing significantly less flux through the boundaries during the initial phases and then maintaining a strong mean field in steady state.

### v3

@@ 3 Theory & Methods
Here $\bm{B}$ is the magnetic field, $\bm{U}=-q\Omega x\bm{\hat{y}}+\bm{u}$ is the total flow velocity (including the background shear $-q\Omega x\bm{\hat{y}}$), $\rho$ is the mass density, $\Omega$ is the local rotation rate, which is the natural time unit for the system, $q\equiv-d\ln\Omega/d\ln R=3/2$ for a Keplerian disk, and $c_{s}$ is the (constant) sound speed.
Although idealized, many studies have shown that the turbulent dynamics induced by Eqs. 1a to 1c are widely varied and dynamically complex, involving the natural generation of long-time coherent large-scale fields and flows, and other complex features. The system also has various undesirable properties such as solutions that depend on the vertical boundary conditions, resolution, and box size (e.g., Simon et al. 2012; Shi et al. 2016; Ryan et al. 2017), as well as the obvious neglect of potentially important global effects (e.g., field-line curvature). For our purposes, it retains two important advantages compared to global simulations: first, it is cheaper computationally than global simulations for similar disk resolutions; second, it provides a much simpler and more controllable testbed for understanding the key physics at play in the system, particularly in steady state. We will see that vertical stratification is absolutely crucial to the physics we describe, meaning that the yet simpler unstratified version of the shearing box would be unsuitable.
@@ 3.1 Static vertical equilibria in the local shearing box
Here $P_{\rm th}\equiv c_{s}^{2}\rho$ is the thermal pressure, $P_{B}\equiv B^{2}/8\pi$ is the magnetic pressure, $\beta(z)\equiv P_{\rm th}/P_{B}$, and $H_{\rm th}^{2}\equiv 2c_{s}^{2}/\Omega^{2}$ is the thermal scale height. We will find that the azimuthal field dominates strongly over the other components. This implies that so long as the turbulence is modestly subsonic where $\beta\gtrsim 1$ and modestly sub-Alfvénic where $\beta\lesssim 1$, then Eq. 2 establishes the relationship between $\rho$ and $B_{y}\approx\sqrt{8\pi P_{B}}$. We will also see that the mean vertical outflows that develop do not provide significant contributions to the overall pressure balance, further justifying the magneto-static assumption.
@@
Figure 1: Various magneto-static equilibrium solutions to Eq. 2, showing $P_{B}$ (solid lines) and $P_{\rm th}$ (dashed lines). Colored line pairs show the Lorentzian-like equilibrium (5) with $a=3$; the blue curves show $H/H_{\rm th}=5$, such that $\beta(z=0)\approx 0.1$, and the orange curves show $H/H_{\rm th}=1.5$, which is just below where $P_{B}$ develops a midplane dip and $\beta(z=0)>1$ (see text). The dotted black line shows the thermal density profile $\exp(-z^{2}/H_{\rm th}^{2})$ and the thin gray curves show the constant-$\beta$ equilibrium (3) with $\beta=0.1$ ($H_{\beta}\approx 3.3H_{\rm th}$), as used to initialize most simulations. All profiles are normalized with $P_{\rm th}(z=0)=1$ for ease of comparison.
@@ 3.2 Boundary conditions
In all simulations we use periodic boundary conditions in $y$, and “shear-periodic” boundary conditions in $x$, whereby quantities are periodic in the frame that remains stationary on the background shear flow $-q\Omega x\bm{\hat{y}}$. However, previous works have shown significant dependence of results on the vertical boundary conditions (e.g., Johansen & Levin, 2008; Gressel, 2013; Salvesen et al., 2016a; Coleman et al., 2017), so we use a number of different choices through this work. All boundary conditions set $u_{x}$, $u_{y}$, and $B_{z}$ to have zero gradient at the vertical boundary, and set the vertical flow $u_{z}$ to have zero gradient unless there is an inflow ($u_{z}z<0$), in which case it is set to $u_{z}=0$ (Stone et al., 1996). The conditions on $\rho$, $B_{x}$, and $B_{y}$ are as follows, with the naming scheme used to label simulations in table 1:
@@
Zero-gradient boundary conditions for $B_{x}$ and $B_{y}$, while $\rho$ is extrapolated into the ghost cells assuming the initial hydrostatic equilibrium as $\rho(z)=\rho_{e}\exp(-z^{2}/H_{\beta}^{2})/\exp(-z_{e}^{2}/H_{\beta}^{2})$ (where the $z_{e}$ and $\rho_{e}$ here refer to the values at the boundary). Following Simon et al. (2011), these boundary conditions have been widely used and allow the total toroidal and radial magnetic flux within the system to vary.
@@
Following Simon et al. (2013) and Salvesen et al. (2016a), these are the same as Outflow, except that they also extrapolate $B_{x}$ and $B_{y}$ similarly to the density, with $B_{i}(z)=B_{i,e}\exp(-z^{2}/H_{\beta}^{2})/\exp(-z_{e}^{2}/H_{\beta}^{2})$. This enhances the removal of magnetic flux from the domain.
@@
Outflow, Simon13, and Power-law effectively differ only in the gradient of $\bm{B}_{\perp}$ and $\rho$ assigned at the boundary, as set by the various extrapolation techniques. For context, the Simon13 hydrostatic scheme (Gaussian extrapolation) is most extreme, fixing quantities to larger negative gradients; the power-law scheme is more moderate, while outflow conditions (with zero gradient) are the least extreme. From a close examination of the shape of $\rho$ near the boundaries with Outflow or Simon13, it is clear that the latter forces the profile to be steeper than is natural. This was the primary motivation for the new Power-law scheme — to disturb the solution as little as possible from the profiles that are naturally set up in coronal regions above the midplane.
@@ 3.3 Initial conditions and numerical setup
Following Salvesen et al. (2016a, b), most of our simulations are initialized with the magnetized constant-$\beta$ equilibrium (3). Most simulations use a vertical domain size $z\in(-L_{z}/2,L_{z}/2)$ with $H_{\beta}^{2}=2$ and $L_{z}=10$, giving significant space away from the midplane to form a low-$\beta$ corona (it transpires that this is necessary to maintain the low-$\beta$ state). Based on the results of Simon et al. (2013) and subsequent works, we use $L_{x}=L_{z}$ and $L_{y}=2L_{z}$ for most runs, but also test the effect of increasing and decreasing both $L_{x}$ and $L_{y}$ by a factor of two (App. A). The equilibrium is then defined by its initial $\beta=\beta_{y0}=8\pi P_{\rm th}/B_{y}^{2}$, which sets the simulation’s isothermal sound speed as $c_{s}^{2}=\Omega^{2}H_{\beta}^{2}(1+\beta^{-1})^{-1}$. In table 1, we label simulations by this initial $\beta$, although the midplane $\beta$ invariably increases significantly from this initial value as the field redistributes and flux is lost through the boundaries.
@@
Nearly all simulations use the piecewise parabolic method with the Van-Leer time integrator and the basic HLL Reimman solver. Although this solver is more diffusive than the HLLD solver, we have found this low-$\beta$ turbulence to be rather sensitive to numerical instabilities, also requiring very conservative choices for the CFL timestep criterion. The strongly magnetized simulations are also quite expensive due to the small timesteps required to capture the fast Alfvén speeds in the very low-$\beta$ coronae. To help mitigate this, which is made worse by the large density variation in such regions, in most simulations we also apply a density floor $\rho_{\rm flr}=10^{-4}\rho_{0}$, as used in previous works [^2] Other approaches include limiting the Alfvén speed (Johansen & Levin, 2008) or a hyper-resistive-like term applied to regions of large |×$|\nabla\times\bm{B}|/\sqrt{\rho}$ (Bai & Stone, 2013). (e.g., Salvesen et al. 2016a; here $\rho_{0}$ is initial density maximum). Unfortunately, this density floor was found to have adversely affected some simulations by limiting the mean density profile far from the midplane. The issue occurs in the high-$\beta$ states at lower $H_{\rm th}$ (see below), which are not our main focus in this work, because the thermal density profile involves an extremely sharp drop off away from the midplane (unlike the low-$\beta$ states, which develop a much flatter power-law profile). In App. C we explicitly test the effect of changing the density floor on the low-$\beta$ state, observing only tiny changes to mean profiles or transport rates.
In all simulations except those with net vertical flux, the mass loss over the duration of the simulation is minimal. In the strongest NVF case with $\beta_{z}=100$, we implement a mass replenishment by multiplying the density by a factor chosen to maintain constant total mass at every time step (as done in previous work; Bai & Stone, 2013).
Note that, compared to previous works, we have run most simulations for a relatively short time (in units of $\Omega^{-1}$). This is motivated by a desire to redeploy computational resources to explore finer resolutions (as needed for $\beta\ll 1$ and small $H_{\rm th}$) and different simulations, rather than measuring precise values of $\alpha$ via long-time statistics. In support of this choice, we note that turbulence in stratified shearing boxes does not yet seem to be converged even at exceptionally fine resolution (Ryan et al., 2017), and detailed study along such lines is beyond the scope of this work. To ensure that the system is truly converged in time, we have run one simulation of the low-$\beta$ state to beyond $100$ orbits; this revealed a nearly constant statistical steady state with only small changes over long timescales.
@@ 3.4 Notation & Units
The shearing-box system is normalized to the arbitrary lengthscale $L_{0}$, yielding freedom to define the simulation’s units. To more clearly compare cases with strong magnetization, we do not normalize our box lengthscales to the thermal scale height $H_{\rm th}$, unlike most previous works. Instead, in most simulations, because $H_{\beta}^{2}=2$ and $\beta<1$, we have $H_{\rm th}<1$ in box units. However, everything in the system could, if desired, be rescaled to $H_{\rm th}$, in which case our domain sizes would vary significantly between simulations and be much larger than $(L_{x},L_{y},L_{z})=(10,20,10)$ (this would also rescale the density and velocity). Indeed, we find that small $H_{\rm th}$ (in box units) seems to be required to see the low-$\beta$ state, which is equivalent to saying that very large domains $L_{z}\gtrsim 20$ are needed in thermal units. All plots and diagnostics are normalized to be dimensionless so can be meaningfully compared to previous works, showing good order-of-magnitude agreement where appropriate (see table 1).
@@ 4 The Parker and magnetorotational instabilities
Here we provide an overview of the magnetorotational and Parker instabilities (Parker, 1958; Balbus & Hawley, 1991), which appear to be the primary linear instabilities driving the dynamics. While these instabilities and their relation to each other have been extensively explored in the literature (e.g., Foglizzo & Tagger, 1995; Balbus & Hawley, 1998) we focus here on the less-well-studied regime with $\beta\lesssim 1$ trans- or super-thermal magnetic fields, so it is helpful to collect and simplify some past results for more straightforward application to our simulations.
@@ 4.1 Magnetorotational instability
Also worth mentioning are other related, but fundamentally global instabilities at $\beta\ll 1$, which were first characterized by Pessah & Psaltis (2005) (the relevant dispersion relation was derived but left unexplored in Chandrasekhar 1961; see also Knobloch 1992; Gammie & Balbus 1994; Das et al. 2018). These instabilities are current driven, depending on the radial gradient of the (super-thermal) toroidal field,
@@ 4.2 Parker instability
The Parker instability (Parker, 1958) is driven by the vertical stratification of the magnetic field. Most commonly, it is studied while ignoring the effect of the radial shear; indeed, Shu (1974) argued that the instability cannot ever be fully stabilized by shear or rotation, because the small-scale radial modes remain unstable (although the growth rates of longer-wavelength modes can be suppressed and shear can play an important role by changing the radial wavenumber in time; Foglizzo & Tagger, 1994, 1995). Convenient expressions are found in Goedbloed & Poedts (2004) Chapter 7, who consider a gravitating slab threaded by a perpendicular magnetic field $B_{y}(z)$ (without shear or rotation). Because the background equilibrium and gravitational acceleration $\hat{g}=-\Omega^{2}z$ are $z$ dependent, a correct treatment requires solving the boundary-value differential equation in $z$ (Horiuchi et al., 1988); however, useful insight can be gained by considering local modes that vary rapidly over the length scale $L\sim z\sim B_{y}/B_{y}^{\prime}$ on which the background varies (i.e., $\delta f\sim\exp(i\kappa z)$ with $\kappa L\gg 1$). This yields an effective local dispersion relation for $\omega(k_{x},k_{y},\kappa)$, which shows that two different types of modes can dominate the instability in different regimes: (i) the “pure interchange mode” with $k_{\|}=k_{y}=0$, which does not bend the magnetic field, and (ii) the “quasi-interchange mode,” with $k_{x}\gg k_{y}$ but $k_{y}\neq 0$, which bends the field. We obtain simple expressions below by simplifying those of Goedbloed & Poedts (2004) assuming an isothermal equation of state, a $\beta\ll 1$ azimuthally directed mean field, and linear vertical gravity $\hat{g}=-\Omega^{2}z$ [^4] In the notation of Goedbloed & Poedts (2004), the disk atmosphere of interest becomes $\Gamma=-\hat{g}\,d\ln\rho/dz$, $\Gamma_{B}=c_{s}^{-2}\hat{g}^{2}$, $\Gamma_{0}=c_{s}^{-2}\hat{g}^{2}/(1+2/\beta)^{2}$, with $X^{2}\equiv\Gamma/\Gamma_{B}$. The quasi-interchange dominates for $\Gamma_{0}\leq\Gamma<\Gamma_{B}$ and the pure interchange dominates for $\Gamma\leq\Gamma_{0}$. In a true global disk, the vertical gravitational force and rotation change with height far from the midplane; given our neglect of rotation and local treatment, the resultant modifications to the Parker instability can be captured by rescaling $\hat{g}$ (Horiuchi et al., 1988).
@@
Figure 2: Space-time ‘butterfly’ plots for a range of ZNVF simulations with different initial conditions and sound speeds. Each case shows $\overline{B}_{y}(z,t)$ in the top subpanel and $\overline{B}_{x}(z,t)$ in the bottom subpanel. In the top row, we see cases that transition into the high-$\beta$ weakly accreting state ($\beta$1, $\beta$0.3, and $\beta$0.1) because they do not retain sufficient magnetic flux during their initial relaxation phase; these develop into well-studied dynamo cycles. The bottom panels show various cases that maintain the lower-$\beta$ strongly accreting state ($\beta$R0.1, $\beta$0.05, and $\beta$0.01); these maintain a quasi-steady state with $\beta\lesssim 1$ in the midplane, losing significantly less flux through the boundaries during the initial phases and then maintaining a strong mean field in steady state.